Circular No. 8711

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2006cg AND 2006ch Two apparent supernovae have been reported from unfiltered CCD

frames: 2006cg by R. Quimby and P. Mondol (in poor weather conditions; cf. IAUC 8622), and 2006ch by K. Itagaki (Teppo-cho, Yamagata, Japan, 0.60-m f/5.7 reflector; communicated by S. Nakano, Sumoto, Japan).

SN  2006 UT  α_{2000}  δ_{2000}  Mag.  Offset
2006cg May 6.23 13^{h}05^{m}02.^{s}32 +28^{o}44′24′′.0 16.6 2′′ W, 3′′ N
2006ch May 9.76 23 47 06.12 +29 28 50.6 16.5 17′′ E, 10′′ S

Additional approximate magnitudes for 2006cg: May 12.25 UT, 16.7; 14.23, 16.5. The estimated positional uncertainty of 2006cg is 1′′.5 in each coordinate. SN 2006cg was found by subtracting a co-addition of images taken between 2004 Dec. 15 and 2005 June 29 (limiting mag ∼ 19.4) from the 2006 May images. A spectrum (range 420–890 nm) of 2006cg obtained on May 17.28 shows it to be a type-Ia supernova around a week past maximum (cf. CBET 511). Additional magnitudes for 2006ch in NGC 7753: 2005 Aug. 30, [19.0; Nov. 10, [19.0; 2006 May 11.764, 16.6; 14.752, 16.8. SN 2006A (cf. IAUC 8656) appeared on the other side of the center of NGC 7753, but no spectroscopic information has been reported on that object.

(99942) APOPHIS L. A. M. Benner, J. D. Giorgini, and S. J. Ostro, Jet Propulsion Lab-

oratory; M. C. Nolan, Arecibo Observatory; and M. W. Busch, California Institute of Technology, report that Arecibo (2380-MHz, 12.6-cm) radar observations of minor planet (99942) = 2004 MN_{4} during May 6.497–6.562 UTC yielded a 5.5σ continuous-wave detection and a Doppler measurement of −118256.8 Hz at May 6.534, for a correction of +0.1 ± 0.1 Hz (+6 ± 6 mm/s) relative to the nominal prediction. An orbit estimation incorporating the new Doppler measurement with 779 optical measurements spanning 2004 Mar. 15–2006 Mar. 26, along with the four Doppler and two range measurements from observations in 2005, increases the 2029 Apr. 13.9 earthcenter miss-distance by 450 km, from 5.86 ± 0.11 to 5.93 ± 0.09 R_{⊕}, and reduces the along-track-position uncertainty at closest approach from ± 730 to ± 570 km (cf. IAUC 8593). The volume of the one-standard-deviation spatial uncertainty region decreases by 23 percent, from 261000 to 201000 km^{3}, and the nominal predicted earth-close-approach distance in 2036 increases from 0.168 to 0.276 AU, moving the statistical earth encounter to a lower-probability region within the distribution of possible orbits.

2006 May 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green